# BadHMP: Backdoor Attack against Human Motion Prediction

### v1

@@ V-B4 Robustness
Most existing backdoor defenses [41, 42, 43, 46] designed for image classification tasks are not applicable to human motion prediction due to the differences in data format. We evaluate the robustness of our attack against fine-tuning [47], which is a general backdoor defense mechanism. The defender is assumed to have a small proportion (30% in this test) of clean training samples to fine-tune the victim model for a few epochs. The results in Table V indicate that although fine-tuning increases the BDE of the victim model slightly, the embedded backdoor can still be successfully activated.
VI CONCLUSIONS
This paper proposed BadHMP, the first backdoor attack targeting human motion prediction task. Two types of triggers and two types of targets are designed to generate smooth and natural poisoned samples. We also introduced a sample selection strategy to further improve the stealthiness of poisoned training samples, making them harder to be detected by the model trainer. The prediction fidelity of the poisoned model to benign input sequences, the activation success rate of target sequences, and the smoothness and naturalness of the trigger sequences of BadHMP are comprehensively evaluated by objective quantitative metrics on two datasets and two model architectures to validate its compliance with the fidelity, effectiveness and stealthiness criteria.

### v2

@@ V-B4 Robustness
Most existing backdoor defenses [42, 43, 44, 47] designed for image classification tasks are not applicable to human motion prediction due to the differences in data format. To evaluate the robustness of our attack, we test its resilience against fine-tuning—a universal defense mechanism. The defender is assumed to retain 30% of the original clean training samples and fine-tunes the victim model for 30 epochs. As shown in Table VI, although fine-tuning increases the BDE, the value remains significantly lower than the corresponding CDE. This demonstrates that the embedded backdoor retains its activation capability, confirming the robustness of our attack against fine-tuning defense.
VI CONCLUSION
This paper proposed BadHMP, a novel backdoor attack targeting human motion prediction tasks. Our key innovation lies in a two-stage poisoning strategy: 1) We extract the motion of a selected limb from a source sample and graft it onto historical input sequences of clean samples. 2) For future sequences, we globally transfer the joint trajectories from the source sample to clean samples. The poisoned samples exhibit provable stealthiness, as both trigger and target are derived from real-world data, inherently satisfying naturalness and accessibility criteria for human bodies. The prediction fidelity of the poisoned model to benign input sequences, the activation success rate of target sequences, and the smoothness and naturalness of the trigger sequences of BadHMP have been comprehensively evaluated by objective quantitative metrics on two datasets and two model architectures.